• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Lindgren

    Log entry time 16:05:12 on December13,2007

    Entry number 220438

    keyword=Day Shift Summary

    Day Shift Summary
    SL/TO: R.A. Lindgren/Konrad Aniol, 3rd: Mustafa Canan

    8:00 left positron, right electron production at 120 degree, E0=739 MeV
    Left p=175.47 MeV, right p=121.17 MeV (unsettled)
    right momentum settled to 1x10-3 by ~9:00 AM.
    8:55 Changed Left Momentum to 163.60, settled at~ 9:00
    10:35 Recycle Left Q2/Q3 and then change Momentum Right to 329.549
    Recycle Right Q2/Q3, Change momentum right 148.021
    12:15 Change momentum Left HRS to P=133.924 MeV


    -------------------------------------------------------
    Left spectrometer, positron at 120 degree, E0=739 MeV

    p=175.47 MeV
    3847, 50 uA on LH2 Start 8:06, 510K
    3848, 50 uA on Left 12C,Start 8:21, 847K
    3849, 50 uA on Left Fe Start 8:40 836K

    p=163.60 MeV
    3850, 50 ua on Left Fe Start 9:00, 236K
    3851, 50 uA on 12C,Start 9:25 250K
    3852 junk
    3853 50 uA on LH2, Start 9:56 500K
    3854 50 uA on Pb, Start 10:08 786K

    P=148.021 MeV
    3855 50 ua on 12C, Start 11:16 138K
    3856 50 uA on Pb, Start 11:38 530K
    3857 50 uA on LH2 Start 11:53 348K
    3858 50 uA on Fe Start 12:09 180K

    P=133.924 MeV
    3859 50 ua 12C Start 12:27 118K

    --------------------------------------------------------
    Right spectrometer, electron production at 120 degree, E0=739 MeV

    P=121.17 Mev (settled at the level of 1x10-3 at ~9:00)
    21837 50 uA on left Fe,Start 9:00 323K
    21838 50 uA on 12C,Start 9:25 300K
    21839 junk
    21840 50 uA on LH2 Start 9:56 160K
    21841 50 uA on Pb Start 10:08 729K

    P=329.549 MeV
    21842 50 uA on LH2 Start 11:53 290K
    21843 50 uA on Fe Start 12:09 82K
    21844 50 uA on 12C Start 12:25 78K

    .
    Stopped beam for beam studies and Hall A access to change polarity and move HRS's
    HRS Left polarity changed and spectrometers mover to 90 degrees. Beam permit establisd at about 2:30
    Since that time there has been no beam due to a problem in the North Linac area.

    1555 Beam back at 50 uA.
    Start the next Run
    Left HRS kinematics 17.5
    Right HRS kinematics 18
    Turning it over to the Swing shift.